### v1

Abstract
As autonomous systems are increasingly deployed in open and uncertain settings, there is a growing need for trustworthy world models that can reliably predict future high-dimensional observations. The learned latent representations in world models lack direct mapping to meaningful physical quantities and dynamics, limiting their utility and interpretability in downstream planning, control, and safety verification. In this paper, we argue for a fundamental shift from physically informed to physically interpretable world models — and crystallize four principles that leverage symbolic knowledge to achieve these ends: (1) structuring latent spaces according to the physical intent of variables, (2) learning aligned invariant and equivariant representations of the physical world, (3) adapting training to the varied granularity of supervision signals, and (4) partitioning generative outputs to support scalability and verifiability. We experimentally demonstrate the value of each principle on two benchmarks. This paper opens several intriguing research directions to achieve and capitalize on full physical interpretability in world models.
keywords:
[…]
1 Introduction
Autonomous systems are increasingly deployed in open and uncertain environments [Saidi et al.(2022)Saidi, Ziegenbein, Deshmukh, and Ernst, Topcu et al.(2020)Topcu, Bliss, Cooke, Cummings, Llorens, Shrobe, and Zuck] and use high-dimensional observations like high-resolution images and LiDAR scans to perceive these environments in necessary detail. To achieve high performance, their planning and control are often implemented with deep learning methods like reinforcement learning (RL) [Yang et al.(2022)Yang, Zhang, Luu, Ha, Tan, and Yu, Garg et al.(2019)Garg, Chiang, Sugaya, Faust, and Tapia]. Since RL training is sample-inefficient, it is impractical to perform in the real world — leading the controllers to be trained “in the imagination” of world models [Ha and Schmidhuber(2018b), Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg].
World models learn to approximate the physical world by predicting future observations based on current observations and actions. Popular neural world models rely on compressing high-dimensional observations into the latent space using an autoencoder. Then these latent values are propagated forward in time based on learned temporal dependencies [Deng et al.(2023)Deng, Park, and Ahn] and decoded into predicted observations. World models can be drastically improved by injecting symbolic physical knowledge into their structure and training process. For example, the approach by [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] automatically discovers physically meaningful variables from raw observations for more stable long-horizon predictions than methods with standard high-dimensional autoencoders. In another instance, [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] decompose control into neuro-symbolic predicates powered by vision-language models and predefined control primitives, thereby improving generalization.
A major challenge of modern world models is their lack of physical interpretability: usually the latent state cannot be easily mapped to physical quantities (e.g., pose or velocity). This limits their use in classical model-based autonomy and the design of physically-grounded rewards for RL. We also cannot obtain physical guarantees from reachability analysis based on world models [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer]. The core reason for this uninterpretability is that deep learning thrives on distributed representations, in which each feature is partially encoded in multiple latent variables [Hinton(1986)]. This challenge is further complicated by partial online observability of the physical state and the difficulty of precisely labeling the data (e.g., indicating which state is riskier in a video).
This paper calls for a paradigm shift from physically informed world models to physically interpretable ones. The former use symbolic physical knowledge to make learning more effective, efficient, and generalizable. The latter creates neuro-symbolic latent representations with explicit physical meaning, thus subsuming physically informed approaches. Physically meaningful representations bring in a plethora of desirable qualities such as reliability, verifiability, and debuggability.
By carefully analyzing the existing world model literature, this paper advances four guiding principles that underlie physical interpretability of learned world representations. Specifically, each principle asserts that physically interpretable world models should:
Principle 1: …be structured according to the physical intent of latent variables.
Principle 2: …learn aligned invariant and equivariant representations of the physical world.
Principle 3: …adapt their training to the varied granularity of supervision signals.
Principle 4: …partition their generative outputs to support scalability and verifiability.
The next section identifies the interpretability gaps in the existing world models, while Section 3 details the four principles. In Section 4, we perform lightweight validation to demonstrate the value of these principles. Finally, Section 5 discusses the newly opened directions for future research.
2 World Models: State of the Art
Foundations of world models. Modern world models have led to state-of-the-art performance in autonomous planning and control while addressing the data-efficiency concerns of standard RL [Deng et al.(2023)Deng, Park, and Ahn, Micheli et al.(2023b)Micheli, Alonso, and Fleuret, Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling]. Modern deep-world models were first proposed by [Ha and Schmidhuber(2018a)]: a variational autoencoder (VAE) generated image representations and a recurrent neural network made latent-space predictions. Later work refined both the encoder-decoder architecture and the surrogate dynamics: PlaNet introduced a recurrent state-space model (RSSM) for prediction [Hafner et al.(2018)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson]; Dreamer propagated gradients back through the imagined trajectories to refine latent prediction [Hafner et al.(2019b)Hafner, Lillicrap, Ba, and Norouzi]; and DreamerV2 extended the RSSM to categorical latent variables [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba]. More recent research integrates autoregressive transformers with self-attention layers to capture detailed temporal dependencies [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling], or diffusion models to mitigate compounding errors [Ding et al.(2024)Ding, Zhang, Tian, and Zheng]. World models have also found success in optimizing planning algorithms for autonomous vehicles in realistic environments: DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] leverages multi-modal inputs to generate realistic video trajectories for policy optimization; DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai], OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu], UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai], and RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] forecast detailed 3D occupancy to inform motion planning and control.
Towards interpretable world models. Despite the strong performance of world models, their interpretability remains a major challenge in most frameworks. Early efforts toward disentangling latent variables (i.e., reducing their mutual dependency) include -VAEs [Higgins et al.(2016)Higgins, Matthey, Pal, Burgess, Glorot, Botvinick, Mohamed, and Lerchner] and causal VAEs [Yang et al.(2021)Yang, Liu, Chen, Shen, Hao, and Wang]. This disentanglement strategy is also employed in driving prediction frameworks like GNeVA [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu] and ISAP [Itkina and Kochenderfer(2022)]. Under the umbrella of world models, G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] investigated a principled modeling framework that inherits interpretable object and context latent separation from various spatial attention approaches [Kosiorek et al.(2018)Kosiorek, Kim, Posner, and Teh, Kossen et al.(2020)Kossen, Stelzner, Hussing, Voelcker, and Kersting, Jiang et al.(2020)Jiang, Janghorbani, Melo, and Ahn, Crawford and Pineau(2020)]. Interpretability can also manifest in the world model’s forward dynamics: [Cui et al.(2020)Cui, Nguyen, Chou, Lin, Schneider, Bradley, and Djuric] propose embedding known dynamics into neural pipelines, while more recent work imposes further physical constraints on system identification [Sridhar et al.(2023)Sridhar, Dutta, Weimer, and Lee], motion prediction [Tumu et al.(2023)Tumu, Lindemann, Nghiem, and Mangharam], and ordinary differential equations with learnable parameters [Linial et al.(2021)Linial, Ravid, Eytan, and Shalit, Zhong and Meidani(2023), Mao et al.(2025)Mao, Gu, Sha, Shao, Wang, and Abdelzaher]. Incorporating partial knowledge of physics with weak supervision has recently improved both the state and dynamics interpretability [Mao and Ruchkin(2024)]. A recent Nature article leveraged the biological alignment of latent representations to predict microbiome community interactions and antibiotic resistance [Baig et al.(2023)Baig, Ma, Xu, and You]. Neuro-symbolic world models have also begun to emerge: VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] learns a set of abstract states and high-level actions for strong out-of-distribution generalization, whereas WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] learns symbolic rules to adapt the dynamics to open world novelty.
Knowledge gap. We observe the lack of world model architectures supporting full physical interpretability, as illustrated in Figure 1 (the underlying literature list can be found in Table 2 in the appendix). Some existing neuro-symbolic architectures scrape the threshold of physically interpretable dynamics, yet lack fluid state representations. On the other hand, powerful multimodal transformer-based architectures preserve the physical context through 3D occupancy, but their prediction mechanisms are black-box. Bridging this gap is key to transitioning from merely physically informed world models to fully physically interpretable ones. Our recent work highlights the need to address these open questions [Lu et al.(2024b)Lu, Afridi, Kang, Ruchkin, and Zheng] and the promise of predictive world models [Mao et al.(2024b)Mao, Sobolewski, and Ruchkin] and their foundation-model variants [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin].
Benefits of Physical Interpretability. Aligning world models with fundamental physical principles (e.g., kinematics and conservation laws) has been shown to improve their out-of-distribution generalization and robustness [Mao and Ruchkin(2024), Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn, Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis, Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl]. These principles prevent latching onto spurious correlations in training and constrain the models to traverse a physically meaningful manifold when extrapolating observations. Going further, physically interpretable representations would lead us to a qualitatively new level of safety and trustworthiness. It would make world models more transparent and debuggable by cross-checking them with real-world physics. It would also make generative components suitable for closed-loop verification of physical properties. Finally, physical representation would drastically improve RL sample efficiency by shrinking the search space to physically feasible solutions.
3 Physical Interpretability Principles for World Models
3.1 Principle 1: Structuring the Latent Space with Prior Knowledge
Although many facets of human cognition remain a mystery, extensive research indicates that distinct regions of the brain are dedicated to unique functions [Luria(1970)]. For example, early studies of patients with brain injuries revealed the existence of functional systems partially dedicated to meaningful tasks such as spatial awareness, voluntary movement, and sensory processing [Luria(1970)]. While modeling the minutia of human cognition is beyond the scope of this principle, this idea of functional organization inspires how world model architectures are increasingly structured.
[…]
Recent work in the context of Branch 1 aligned latent representations with physical properties [Mao and Ruchkin(2024)]. Although modeling agent interactions is a well-explored topic, early work by [Kipf et al.(2018)Kipf, Fetaya, Wang, Welling, and Zemel] revealed how physical interactions between agents can be learned through graph neural networks (GNN) in an unsupervised manner. Later advances in learning physical interactions within the context of world models by [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] proposed a method that constructs a separate latent representation using a GNN to capture agent occlusion and interaction relationships. Finally, Branch 3 follows the typical strategy for creating uninterpretable world models and is considered a useful layer to the structured latent space [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn].
Latent space structuring has become increasingly prevalent in improving the performance of planning and control. To this end, [Lu et al.(2024a)Lu, Zhan, Tomizuka, and Hu] developed a goal-based neural variational agent (GNeVA) that uses separate polyline embeddings for the agent and the map to achieve interpretable generative motion prediction. Similarly, [Itkina and Kochenderfer(2022)] designed an interpretable car trajectory prediction that integrates three distinct workflow branches: agent states, high-definition maps, and social context.
3.2 Principle 2: Exploiting Invariances and Equivariances in Input and Latent Spaces
Deep neural networks have achieved impressive performance due in part to their ability to learn rich distributed representations from training data. Rather than simply memorizing examples, these models construct hierarchical feature embeddings that capture patterns in the data to generalize to i.i.d. samples [Hinton(1986)]. Nevertheless, training a model to internalize and imagine the world in a human-like manner from scratch is still a nontrivial challenge at best ([Ha and Schmidhuber(2018a)]). Encoding high-dimensional observations (e.g., images) through commonplace embedding methods (e.g., through autoencoders or encoder-only transformers) leaves the latent representation generally uninterpretable and task-agnostic. This raises concerns about whether the latent space is distorted by spurious correlations or if it effectively encodes the details necessary for discriminating between features that should remain functionally disentangled.
Invariance and equivariance relations are one way to address uninterpretability in representation learning models. These terms are often used to classify representations based on their response to observation space transformations. If the representation of shifts in an expected manner due to a transformation , then the representation model is said to be equivariant to that transformation. Likewise, if the representation does not shift under the transformation, then the model is said to be invariant to the transformation.
To this end, [Zhang et al.(2020)Zhang, Lyle, Sodhani, Filos, Kwiatkowska, Pineau, Gal, and Precup] utilize bisimulation metrics to learn latent representations of roadway obstacles that are invariant to obstacle type, size, and brightness. The dual of avoiding explicitly irrelevant details is explicitly learning relevant details. [Pol et al.(2020)Pol, Kipf, Oliehoek, and Welling] introduce contrastive loss to enforce action-equivariance on learned representations. However, we still lack ways to effectively incorporate these expert conceptual priors into world models. [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters] innovate a symmetric embedding network for world models that learns simple latent space transformations from complex observation space transformations.
We classify representation models along two dimensions: (1) the nature of their transformation response (invariance versus equivariance) and (2) their degree of human alignment (aligned versus misaligned). A representation that is aligned-invariant remains unchanged when an observation undergoes a meaning-preserving transformation, while an aligned-equivariant representation transforms predictably when the observation’s meaning is altered. In contrast, a representation is misaligned-invariant if it does not change under meaningful effects made to the observation (suggesting underfitting), and it is misaligned-equivariant if it changes in response to an observation transformation that should not affect the underlying meaning (suggesting a domain shift). Our training objective is to achieve invariance and equivariance alignment by ensuring that the post-transformation representations accurately reflect our human interpretation of the change.
Definition 3.1 (Equivariant Representation).
Let be a function that maps each observation to a latent representation . Let be a finite set of transformation pairs such that and . We say that is equivariant to if . Invariance is a special case of equivariance where .
Following Definition 3.1, a simple loss function promotes aligned invariance and equivariance:
In practice, deciding which features the representation should be invariant or equivariant to solely depends on the task at hand and the discretion of the expert engineer. In autonomous driving, consider a recovery component that overrides the default controller during inclement weather. This component’s representations of real-time images might change when the sky is cloudy. On the other hand, the world model of the default controller might be invariant to the features of the sky.
3.3 Principle 3: Multi-Level Supervision for Latent Representations
World models must bridge the gap between rich observations and physical meaning. This requires adapting the training strategy based on the type and quality of available supervision signals [Lee et al.(2013), Chen et al.(2020)Chen, Kornblith, Norouzi, and Hinton]. Instead of treating all supervision signals in the same way, effective world models carefully exploit supervision signals — whether they are precise, coarse, or missing — by selecting the most fitting training method to align with the underlying physical system.
Multi-level supervision tailors the loss functions and training process to the granularity of supervision signals. For instance, physical state labels allow for the direct alignment of latent representations with real-world quantities using supervised loss. When such labels are unavailable, temporal consistency and smoothness of trajectories can serve as implicit regularization techniques to constrain learned representations. Finally, self-supervision can leverage data-driven structures to discover meaningful latent representations in entirely unsupervised settings.
Supervised Learning: Direct supervision connects latent values with physical states. In many cases, supervision signals are introduced directly into the embeddings to capture key features from labeled data [Zhuang et al.(2015)Zhuang, Cheng, Luo, Pan, and He]. For example, in low-dimensional systems with position and velosity states , additional latent dimensions () can improve reconstruction quality and stability [Chen et al.(2016)Chen, Duan, Houthooft, Schulman, Sutskever, and Abbeel, Alemi et al.(2018)Alemi, Poole, Fischer, Dillon, Saurous, and Murphy, Rezende et al.(2014)Rezende, Mohamed, and Wierstra].
Semi-Supervised Learning: When the labels are only available for some data, semi-supervised techniques can refine representations. Pseudo-labeling (e.g., Mean Teacher [Tarvainen and Valpola(2017)] and FixMatch [Sohn et al.(2020)Sohn, Berthelot, Carlini, Zhang, Zhang, Raffel, Cubuk, Kurakin, and Li]) utilizes both labeled and unlabeled data to iteratively improve the latent space. In Motion2Vec [Tanwani et al.(2020)Tanwani, Sermanet, Yan, Anand, Phielipp, and Goldberg], a small amount of labeled data is first used to initialize the embedding space; subsequently, RNNs predict pseudo-labels for unlabeled data, allowing the model to iteratively refine both the embedding and segmentation components.
Weak Supervision: Noisy or coarse labels, such as position constraints , can be utilized via the trajectory smoothness loss: . Temporal models like Kalman filters [Kalman(1960)] stabilize noisy trajectories in tasks such as autonomous driving. Interval signals as weak supervision can be directly incorporated into the loss function [Mao and Ruchkin(2024)] or combined with contrastive learning to reinforce constraints [Sorokin and Gurevych(2017)].
Self-Supervised Learning: In the absence of labels, contrastive learning [Chen et al.(2020)Chen, Kornblith, Norouzi, and Hinton] aligns latent representations with task-specific similarity metrics (e.g., Euclidean distance or structural similarity). Contrastive world models [Poudel et al.(2022)Poudel, Pandya, and Cipolla] explicitly employ representation learning losses to map similar states closer in the latent space. Plan2Explore [Sekar et al.(2020)Sekar, Rybkin, Daniilidis, Abbeel, Hafner, and Pathak] generates self-supervised uncertainty-driven objectives to guide the representations.
Combining Supervision Levels: For a given dataset with multiple supervision signals, the training objective must flexibly incorporate relevant terms to ensure that the latent space aligns with the physical world across diverse scenarios. We advocate for using every available supervision opportunity. Given explicit state labels, use direct supervision. When only partial information is available, use weakly supervised constraints to refine the representations. In addition, use self-supervision to extract the sequential physical patterns of observations. Since all these supervisory signals come from the physical world, we expect their multi-level integration to improve physical interpretability.
3.4 Principle 4: Output Space Partitioning for Verifiability
Ensuring the safety of vision-based autonomy is a critical and open challenge [Teeti et al.(2022)Teeti, Khan, Shahbaz, Bradley, Cuzzolin, and De Raedt, Geng et al.(2024)Geng, Dutta, and Ruchkin]. The verification of such systems remains difficult due to high-dimensional image inputs: traditional techniques struggle to handle this complexity, making it essential to develop new principles for pre-deployment safety guarantees [Althoff(2015), Păsăreanu et al.(2023)Păsăreanu, Mangal, Gopinath, Getir Yaman, Imrie, Calinescu, and Yu]. A recent approach employs a generative image model to map the physical state to the observed image, in turn fed into a state estimator or controller [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer, Cai et al.(2024)Cai, Fan, and Bak]. Unfortunately, due to uninterpretable latent states, such “verification modulo generative models” does not provide guarantees regarding the physical world. Furthermore, it does not scale to large image sizes.
We propose to partition the generated image into physically meaningful parts to enable the safety verification of vision-based controllers. Specifically, a world model will contain multiple generators of output signals — each dedicated to its own image region. Each generator would be separately verifiable, and the results would be combined to provide world model-wide guarantees. This principle reduces each generator’s size, making the analysis and execution of world models more parallelizable and scalable. Moreover, when applied to physically interpretable latent states, this principle enables the transfer of verification guarantees to the physical world: the generators will represent the relationship between images and physical states, not uninterpretable latent ones.
Definition 3.2 (Partitioned World Model Generation).
A world model decoder translates a latent state into a generated high-dimensional observation , expressed as , by minimizing the reconstruction error between the original and reconstructed observations. Each image segment is produced by a separate decoder: . The combined generated image is represented as , where is a signal composition operation (e.g., overlaying image segments). The corresponding loss function is:
The question of automatic partitioning of world model outputs can be answered by zero-shot approaches like the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]. Recently, SAM was used to segment images to improve image and safety prediction [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin]. A similar partitioning was used in the action space to scale up the verification of vision-based controllers via multiple low-dimensional approximations [Geng et al.(2024)Geng, Dutta, and Ruchkin]. Principle 4 propagates the physical meaning from different parts of the world model (established in Principle 1) to its generative outputs, effectively linking the high-dimensional observation space with a lower-dimensional state representation.
4 Experimental Validation
The objective of our experiments is to evaluate the impact of the four proposed principles on the interpretability of world model representations. We expect each principle to improve the prediction of future physical states compared to a baseline interpretable world model. The success is measured by the mean squared error (MSE) of state predictions over varying prediction horizons.
[…]
| World model | Environment | Average MSE | Average SSIM | Model Size |
| Baseline (monolithic) | Cart Pole | 0.02856 | 0.997122 | 200,259 |
| Partitioned 3-way | Cart Pole | 0.05176 | 0.995614 | 144,665 |
| Baseline (monolithic) | Lunar Lander | 0.18801 | 0.8686 | 360,773 |
| Partitioned 3-way | Lunar Lander | 0.306 | 0.6289 | 78,101 |
Principle 1: Here we split the encoder into the image part for extracting low-level visual features and the state part that produces values of physical variables. The latent vector size is the same for the baseline and the modified models. Figures 3A and 3C show that Principle 1 significantly reduces the MSE for longer horizons, highlighting the stability that comes from physical interpretability.
Principle 2: For the lunar lander, we add a translation to both the observation and position, ensuring the equivariance to translation. For the cart pole, a translation is applied to the cart and a rotation to the pole, with corresponding changes to the latent state. Figures 3A and 3C show that Principle 2 reduces prediction error across all prediction horizons, confirming the value of equivariance.
Principle 3: Here we train world models in semi- and weakly-supervised settings: (1) only static information (position, angle) is supervised, while dynamic (velocity) is unknown; (2) velocity is estimated from positions/angles, adding supervision through physical knowledge. Figures 3B and 3D show that weak physical supervision improves prediction quality at all prediction horizons.
Principle 4: We partition the original cartpole and lunar lander images into three parts with SAM. Then three smaller decoders are trained to generate each image part, which are combined in the end (as illustrated in Figure 4 in the appendix). The baseline decoder has one linear layer, two convolutional layers, and a 4-dimensional encoded feature map. Our partitioned decoder only contains one linear layer and one smaller convolutional layer. In both the baseline and its partitioned version, the latent space has 4 dimensions for the cartpole and 8 for the lunar lander. The partitioned generator inputs are the exact physical states, while the baseline approach’s uninterpretable latents. Our partitioning reduces 200,259 parameters in the baseline to only 144,665 parameters. Despite that, the reconstruction quality remains comparable, as demonstrated in Table 1.
5 Future Research Directions
A. Extracting Physical Knowledge from Foundation Models. It is difficult for humans to externalize their implicit knowledge of the physical world [Trager et al.(2023)Trager, Perera, Zancato, Achille, Bhatia, and Soatto, Xu et al.(2024)Xu, Liu, Sokolsky, Lee, and Kong]. Having absorbed humanity-scale data patterns, large language models are promising sources of implicit and plausible physical knowledge. We intend to investigate how to extract candidate dynamics templates, invariances, and equivariances. An important step is validating the candidate information (e.g., via open datasets) before incorporating it into the world model training.
B. Physically Aligned Multimodality. Reliable multimodal world models are urgently needed in many autonomous systems [Gupta et al.(2024)Gupta, Gong, Ma, Pawlowski, Hilmkil, Scetbon, Rigter, Famoti, Llorens, Gao, Bauer, Kragic, Schölkopf, and Zhang, Zheng et al.(2025)Zheng, Li, Ruchkin, Piskac, and Pajic]. However, the consistency of predicted modalities has been a challenge for learned representations [Lu et al.(2024b)Lu, Afridi, Kang, Ruchkin, and Zheng]. We suggest the use of physically meaningful representations in making image and LiDAR predictions consistent on real-world datasets such as nuPlan [Caesar et al.(2022)Caesar, Kabzan, Tan, Fong, Wolff, Lang, Fletcher, Beijbom, and Omari] and Waymo Open [Sun et al.(2020)Sun, Kretzschmar, Dotiwalla, Chouard, Patnaik, Tsui, Guo, Zhou, Chai, Caine, Vasudevan, Han, Ngiam, Zhao, Timofeev, Ettinger, Krivokon, Gao, Joshi, Zhang, Shlens, Chen, and Anguelov].
C. Interpretable Uncertainty in World Models. Commonplace uncertainty quantification techniques for deep learning models struggle to express the uncertainty in the terms relevant to the application domain [Gal and Ghahramani(2016), Kendall and Gal(2017)]. Traditional Bayesian approaches and ensemble methods often focus on model uncertainty but fail to capture the structured uncertainty inherent in physical systems [Zhang et al.(2019)Zhang, Lu, Guo, and Karniadakis]. In contrast, uncertainty estimation within physically meaningful latent representations allows for more interpretable and actionable uncertainties. We consider it fruitful to develop an uncertainty quantification method based on distributions over physically meaningful latent states and partitioned outputs, which can facilitate robust decision-making and improve reliability in downstream tasks [Depeweg et al.(2018)Depeweg, Hernandez-Lobato, Doshi-Velez, and Udluft].
D. Unified Training Pipeline. This paper outlined a variety of training objectives and supervision strategies for physically interpretable world models. However, when their combinations are used, the convergence and stability of training remain elusive [Sener and Koltun(2018)]. We recommend developing an automated training pipeline that will combine and tune different losses to ensure reliable training [Li et al.(2018)Li, Jamieson, DeSalvo, Rostamizadeh, and Talwalkar].
E. Integrating World Models into Classical Autonomy. Physically meaningful states enable high-performance components of world models to serve as state estimators, trajectory predictors, and models for verification [Mao and Ruchkin(2024)]. These capabilities allow combinations between two previously incompatible first-principles and end-to-end learning. We intend to enhance classic autonomy tasks with world-model components to improve their performance while preserving their reliability and verifiability.
The authors thank Vedansh Maheshwari, Mrinall Umasudhan, Rohith Reddy Nama, Sukanth Sundaran, and Liam Cade McGlothlin for their experiments with neural representations. This research is supported in part by the NSF Grant CCF-2403616. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation (NSF) or the United States Government.
Appendix
Experimental Details
[…]
Principles 1–3. Our world model employs a VAE for encoding/decoding high-dimensional image observations and an LSTM time-series predictor for modeling state transitions in the latent space. The encoder consists of three convolutional layers with increasing feature maps (16, 32, 64) and ReLU activations, downsampling the input image through strided convolutions. The latent representation is parameterized by two fully connected layers ( and ), each mapping the encoded feature vector to a latent space of 64 dimensions. The decoder reconstructs the input image using a fully connected layer followed by three transposed convolutional layers, producing a three-channel output with a sigmoid activation. The VAE is trained using the Adam optimizer with an initial learning rate of 0.001, incorporating learning rate decay to stabilize convergence.
The input to the LSTM consists of 64-dimensional latent representations extracted by the VAE. The network comprises two LSTM layers with a hidden size of 64, followed by a fully connected output layer mapping to a 64-dimensional output representing the predicted latent state at the next time step. The LSTM predictor is trained using the Adam optimizer with an initial learning rate of 0.001 and also incorporates learning rate decay. The objective is to minimize the MSE between predicted and true latent representations over time.
Principle 4. Our decoder network maps low-dimensional physical state representations to high-dimensional images using a series of transposed convolutional layers. Using a fully connected layer, the decoder first maps the input state (four-dimensional vector in cartpole; eight-dimensional vector in lunar lander) to a high-dimensional feature space. This produces an intermediate representation of size 3×16×24×24. The image output is further refined through independent transposed convolutional layers, each producing a separate image (three independent layers for each segment image for cartpole and lunar lander). The model is trained using the Adam optimizer with an initial learning rate of 0.001. Training is conducted with mini-batches of size 64, incorporating validation loss tracking to ensure generalization. The loss function is a -weighted combination of the reconstruction MSE of the overall reconstructed image and each segmented part.
For the partitioned loss function in Equation 1, the choice of plays a crucial role in image generation behavior: (a) If is too small ), the model fails to separate the three parts, blending “shadows” of the original image into the outputs; (b) If is too big (), the three parts are completely disconnected, leading to inferior reconstruction quality. Through hyperparameter tuning, we found that setting provides an optimal balance between the quality of the separation and the reconstruction in both case studies.
Additional Illustrations
Table 2 lists the literature on world models and indicates the extent of state/dynamics interpretability and adherence to the proposed principles.
Figure 4 shows example observations and their partitioned reconstructions for Principle 4.
Figure 5 shows the imperfect part-wise reconstruction for inadequate values of .
Short Name Reference Principle 1 Principle 2 Principle 3 Principle 4 State interp. Dyn. interp. WM [Ha and Schmidhuber(2018b)] PlaNet [Hafner et al.(2019a)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson] Weak Weak Dreamer [Hafner et al.(2020)Hafner, Lillicrap, Ba, and Norouzi] Weak Weak G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] Strong Weak Weak Moderate Weak AWM [Kim et al.(2020)Kim, Sano, Freitas, Haber, and Yamins] Weak Weak Weak Plan2Explore [Sekar et al.(2020)Sekar, Rybkin, Daniilidis, Abbeel, Hafner, and Pathak] Weak Weak Weak Pathdreamer [Koh et al.(2021)Koh, Lee, Yang, Baldridge, and Anderson] Weak Weak Strong Weak DreamerV2 [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba] Weak Weak Weak NSV [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] Strong Strong DayDreamer [Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg] Weak Weak DreamingV2 [Okada and Taniguchi(2022)] Weak Weak Weak SEN [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters] Strong Moderate STEDI [Nakano et al.(2022)Nakano, Suzuki, and Matsuo] Strong Moderate Strong DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] Weak Strong Weak GAIA-1 [Hu et al.(2023)Hu, Russell, Yeo, Murez, Fedoseev, Kendall, Shotton, and Corrado] Strong Moderate IFactor [Liu et al.(2023)Liu, Xiao, and Belta] Moderate Moderate IRIS [Micheli et al.(2023a)Micheli, Alonso, and Fleuret] Weak Weak MTS3 [Shaj et al.()Shaj, Zadeh, Demir, Douat, and Neumann] Weak Weak Weak Weak Denoised MDP [Wang et al.(2023a)Wang, Du, Torralba, Isola, Zhang, and Tian] Moderate Moderate WM2WM [Wong et al.(2023)Wong, Grand, Lew, Goodman, Mansinghka, Andreas, and Tenenbaum] Strong Moderate MWM [Seo et al.(2023)Seo, Hafner, Liu, Liu, James, Lee, and Abbeel] Weak Weak OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu] Strong Strong Strong RAP [Hao et al.(2023)Hao, Gu, Ma, Hong, Wang, Wang, and Hu] Moderate S4WM [Deng et al.(2023)Deng, Park, and Ahn] Moderate SWIM [Mendonca et al.(2023)Mendonca, Bahl, and Pathak] Moderate Moderate Moderate TWM [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling] Weak Weak UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai] Strong Strong Strong WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] Strong Moderate Strong THICK [Gumbsch et al.(2023)Gumbsch, Sajid, Martius, and Butz] Weak Weak Weak AVID [Rigter et al.(2024)Rigter, Gupta, Hilmkil, and Ma] Strong Moderate CMIL [Kolev et al.(2024)Kolev, Rafailov, Hatch, Wu, and Finn] Strong Weak DreamerV3 [Hafner et al.(2024)Hafner, Pasukonis, Ba, and Lillicrap] Weak Weak Weak Weak DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai] Strong Moderate Weak DWM [Ding et al.(2024)Ding, Zhang, Tian, and Zheng] GaussianWorld [Zuo et al.(2024)Zuo, Zheng, Huang, Zhou, and Lu] Moderate Strong Strong Moderate Genie [Bruce et al.(2024)Bruce, Dennis, Edwards, Parker-Holder, Shi, Hughes, Lai, Mavalankar, Steigerwald, Apps, Aytar, Bechtle, Behbahani, Chan, Heess, Gonzalez, Osindero, Ozair, Reed, Zhang, Zolna, Clune, Freitas, Singh, and Rocktäschel] Moderate Moderate HarmonyWM [Ma et al.(2024)Ma, Wu, Feng, Xiao, Li, Hao, Wang, and Long] Weak Weak OccWM [Zhang et al.(2024)Zhang, Xue, Yan, Zhang, Qiu, Bai, Liu, Cui, and Li] Moderate Strong Strong Weak CovWM [Popov et al.(2024)Popov, Degirmenci, Wehr, Hegde, Oldja, Kamenev, Douillard, Nistér, Muller, Bhargava, Birchfield, and Smolyanskiy] Weak Weak Weak Weak NWM [Bar et al.(2024)Bar, Zhou, Tran, Darrell, and LeCun] OccLLaMA [Wei et al.(2024)Wei, Yuan, Li, Hu, Gan, and Ding] Strong Strong PIWM [Mao and Ruchkin(2024)] Strong Strong Moderate Strong R-AIF [Nguyen et al.(2024)Nguyen, Yang, Buckley, and Ororbia] Weak Weak Weak RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] Strong Strong Strong Think2Drive [Li et al.(2024)Li, Jia, Wang, and Yan] Weak Weak TransDreamer [Chen et al.(2024)Chen, Wu, Yoon, and Ahn] Weak Weak Weak VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] Strong Strong Moderate WorldGPT [Ge et al.(2024)Ge, Huang, Zhou, Li, Wang, Tang, and Zhuang] Moderate Moderate Moderate Moderate Our future vision Strong Strong Strong Strong Strong Strong